# Accounting for the geometry of the lung in respiratory viral infections

### v1

@@ 3.3 Tissue geometry and aspect ratio influence the fate of competing viral lineages
As a starting point, we sought to compare the lineage dynamics on tubes of varying circumferences. We did so by constructing tubes of length 500 cells and different choices of circumference $C$. As in an earlier result, we ran simulations of our model by seeding infection with four initially infected cells randomly placed on the left edge of the tube, with the exception that these cells are now assumed to be infected by different viral lineages. As a control, we also compared each case with a “naive” approach, where we constructed an approximately square tissue with the same number of cells, and applied toroidal periodic boundary conditions. Note that we define a tissue as “square” by having equally many cells in width as in length, however, due to the hexagonal packing of the cells, the physical dimensions of the sheet are rectangular. In the absence of tissue edges in this case, the four seed cells were placed randomly in the toroid. This control case is representative of the standard modelling approach in the literature where the geometry of the lung is not considered. Schematics and simulations on different geometries are presented in Figure 5. We show schematics of the tube and toroid geometries in Figure 5(a)–(b). We compare the results obtained for a narrow tube (with a circumference of 8 cells) to those obtained for a wide tube (circumference of 64 cells) — along with their corresponding toroids — in Figure 5(c)–(j).
@@
In order to quantify these observations, we ran 100 simulations on both the narrow and wide tube and their corresponding toroids and reported the distribution of proportions of the tissue infected by each lineage. We show these results as histograms in Figure 5(i)–(j). Figure 5(i)–(j) shows two key results. Firstly, the distributions for the toroids and tubes are qualitatively completely different. For the tubes, lineages are by far most likely to infect <10% of the tissue, with higher proportions markedly less likely, demonstrating dominance by a single lineage (or a small number of lineages). By contrast, for the toroids, density is centred at 20–30%, corresponding to a relatively even split of the tissue between the four lineages. This distribution is more tightly peaked for the larger toroid due to the larger cell population. This finding suggests that taking the typical approach to multicellular modelling, that is, in taking the cell sheet to be approximately square and with toroidal boundaries, is insufficient for capturing the lineage dynamics seen on more realistic lung tissue geometry. Figure 5(i)–(j) also show a small but not insignificant probability of a single lineage dominating the cell population (>80%) in the narrower tube. This is not observed for the wide tube.
3.4 Branching of lung tissue promotes loss of within-host viral diversity and spatial isolation of viral lineages
Having established the results in the previous section for a single tube, we next sought to explore how viral lineage dynamics were impacted by the inclusion of branching structure in the tissue geometry. To do so, we first explored how the lineage dynamics on the branching tree were impacted based on whether the infection started in the upper or lower branches. We conducted a set of 100 simulations on a branching tree where, as with the tubes, infection was initiated with four infected cells of different lineages, randomly placed on the left, open edge of the tree (left edge of Generation 1). Then, for comparison, we also carried out 100 simulations where the seed cells were randomly placed within a single branch on the right, branching edge of the tree (right edge of Generation 5). These two scenarios are chosen to represent, in an abstract way, infections of the upper and lower respiratory tract, respectively. As in Section 3.3, we again use a branching tree with maximum circumference of 64 cells and five generations of length 100 cells, such that the overall dimensions of the tissue are the same as that of the wide tube, as previously defined. Simulations on tube and branching geometries and analyses are shown in Figure 6.
@@
In Figure 6(e)–(f) we show representative snapshots of the unrolled branching tissue for both cases of the source cell positions, along with the tube snapshots from the previous section for comparison (Figure 6(c)–(d)). Figure 6(b) gives a schematic of these simulations. Figure 6(e)–(f) provide clear visual evidence for the sharp contrast between the lineage dynamics of infection on the branching tree starting in the upper or lower branches. These snapshots also suggest that the tissue is more likely to be dominated by a single lineage in the branching tree case (seeded on the branching edge) compared to the narrow tube case due to the larger overall cell population. As can be seen in Figure 6(f), by the time the infection spreads the length of the tree and reaches Generation 1, there is essentially only one lineage left but the entire upper half of the tree remaining to be infected, resulting in that lineage infecting a far greater overall proportion of the total cell population.
@@
As another descriptive tool for analysing the lineage dynamics of infections on varying tissue geometries, we next developed a notion of lineage “extinction” if a lineage is not found beyond some specified cut-off depth into the tissue (we use a cut-off of 300 cells). We denote by $P_{\text{\tiny ext}}$ the probability that, in a given geometry, a given lineage will go extinct. Figure 6(b) shows a schematic of this construction. Clearly, extinction as we have described it relies on a single direction of viral invasion, and is therefore not well defined for infection of the branching tube from the branched edge. In the narrow tube, we computed $P_{\text{\tiny ext}}=0.633$, meaning that a given lineage was more likely than not to go extinct, whereas for the wide tube, we found $P_{\text{\tiny ext}}=0.245$, meaning extinction only occurred around 25% of the time. Interestingly, on the branching tree (with source on the open edge), the probability of extinction was slightly reduced at $P_{\text{\tiny ext}}=0.200$, despite containing the same number of cells as the wide tube but with the inclusion of narrow branches within its geometry. This finding results from the fact that the branching structure acts to keep lineages separate, which slightly reduces the chance of their being crowded out. This can be seen in the lineage structure at the right end of both the wide tube and the branching tube (with source on the open edge). While in the tube case the remaining lineages are well mixed across the tube’s circumference, in the branching case, each of the narrow terminal branches is almost entirely uniform in lineage. When infection is seeded in these narrow branches, however, the branching structure does not act to separate and protect lineages since, as was observed in the narrow tube (Figure 6(c)), lineages are rapidly crowded out before they have a chance to be separated and lineage diversity is quickly lost.
@@
To further quantify the difference in lineage structure on the branching tree (seeded on the open edge) and the wide tube, we analysed properties of the lineage dynamics along the depth of the tissue. To do so, for a given depth $d$ cells along the tissue, we examined the band of cells with depth $d-b/2$ to $d+b/2$ (where $b$ is the bandwidth, taken to be 10 cells). We sketch this computation process in Figure 7(a). For each band of cells analysed, we computed both the number of lineages found in the band, and a clustering metric, $\kappa_{\text{\tiny lin}}$, as a measure of the extent to which lineages are clustered together. This metric, which is similar to one which we developed elsewhere [30], is computed on a cell population after an infection is complete and simply reports the mean proportion of a cell’s neighbours which are of the same lineage as the cell. We show a schematic of this calculation in Figure Figure 7(b). We plot both the number of lineages and the clustering metric $\kappa_{\text{\tiny lin}}$ as a function of band depth $d$ in Figure 7(c)–(d). In both cases we indicate the mean along with the 10th and 90th percentiles. Figure 7(c)–(d) both show that for the first approximately 200 cells of depth, the two geometries are relatively indistinguishable, and the difference in the final number of lineages remaining is not substantial compared to the noise of the system. However, for $\kappa_{\text{\tiny lin}}$, we observe that, as well as there being a notable divergence in the means for the branching tree and tube case, there is also a substantial difference in the amount of variation in each case. While $\kappa_{\text{\tiny lin}}$ is an extremely noisy metric towards the right end of the tissue, it is very consistently close to 1 for the branching tube. This confirms our observation from the snapshots that while lineages are somewhat mixed in the tube case, the branches in Generation 5 of the branching geometry generally contain only a single lineage. Note that the reason for the small range of $\kappa_{\text{\tiny lin}}$ values (around [0.7,1]) is a result of the high proportion of infections arising from cell–to–cell infection, leading to a high probability of same-lineage neighbour cells. For a different ratio of cell–to–cell to cell-free infections, it is possible that the difference in $\kappa_{\text{\tiny lin}}$ trajectories for the branching tree and tube geometries would be more pronounced.
@@ 3.5 Immune response depends on the location of infection within a branching structure
Our simulations so far have not explicitly considered the role of the immune response. Resolution of infection in our model is controlled only by the availability of the susceptible cell population, a defining feature of target cell–limited models. However, actual respiratory infections in vivo do not result in the destruction of the entire lung epithelium. Instead, only reasonably localised regions are affected, such as the upper airways or the alveolar passages, before the host immune response or drug intervention results in the clearing of infection [15]. We sought to investigate the behaviour of a basic model of immune activity, and how it might interact with infections seeded in different regions of the branching tree geometry.
@@
Figure 8: (a) Schematic of a simple model for the immune response in the lung. We first specify a threshold, $F^{\text{\tiny thresh}}$ for cumulative infected cell proportion. Once the infection meets the threshold, we assume the immune response is triggered, and then takes $t_{\text{act}}$ hours to activate. Once activated, we assume that the immune response immediately clears the infection. We report the cumulative infected proportion of the tissue at this point, $F^{\infty}$, as a measure of damage to the tissue. (b) Time series for the cumulative infected proportion for infections on the branching tree, where infection is initiated with a randomly placed single cell on either the left edge of Gen. 1 (the open edge of the geometry), or on the right edge of Gen. 5 (the branched edge). In each case, we plot ten trajectories and the mean, along with an indicative value of $F^{\text{\tiny thresh}}$. (c)–(d) Cumulative tissue damage on the branching tree after immune action for varying $F^{\text{\tiny thresh}}$ and fixed $t_{\text{\tiny act}}=30$h, where the infection is seeded on the open edge or the branched edge, respectively. We show violin plots for $F^{\text{\tiny thresh}}$ across 100 simulations (mean in black). We also plot in red the predicted final cumulative damage where the immune response is applied to the mean of the trajectories.
@@
As in the previous result, we sought to apply our simple model of the immune response to infections of the branching tree beginning from either the open left edge of Generation 1, or from the highly branched right edge of Generation 5. We recall that these two scenarios are chosen to represent upper– and lower–respiratory infections in vivo, respectively. In each scenario, as in Figure 3, we initiate infection with a single infected cell randomly placed on the specified edge of the tissue. We plot representative behaviour of both infection scenarios in Figure 8(b). Figure 8(b) shows the cumulative infected proportion over time for infections seeded on either the open or the branched edge. In each case we plot ten representative trajectories and the mean trajectory. The dynamics of infections progressing from either end of the tissue geometry are substantially different from each other. For infections initiated on the open edge, the cumulative infected proportion grows roughly linearly throughout, whereas for infections spreading from the branched edge, initial growth is slow and exponential, followed by a change in convexity at a cumulative infected proportion of around 0.7. There is little variation between individual simulations of the model.
@@
We began by testing the final cumulative damage to the tissue, $F^{\infty}$, for each of these infection seed positions for varying values of the detection threshold, $F^{\text{\tiny thresh}}$. We ran 100 simulations of the model for infections of the branching tree seeded either on the open or the branched edge, and applied our immune model for a range of $F^{\text{\tiny thresh}}$ values between 0.1 and 0.5. Here, we have kept $t_{\text{\tiny act}}$ fixed at 30h. We constructed violin plots for the distribution of $F^{\infty}$ values at each $F^{\text{\tiny thresh}}$ value, and annotated these with the mean $F^{\infty}$ value. We compared this to the $F^{\infty}$ computed on the mean trajectory for both source positions, which we also show on the same axes. That is, we compare the mean of the $F^{\infty}$ on the individual trajectories and compare this to the $F^{\infty}$ on the mean trajectory. We show these results for both the open and branched edge source in Figure 8(c) and (d), respectively. Figure 8(c)–(d) show that cumulative damage increases roughly linearly on this range of detection thresholds for the open edge case, while for the branched edge case, damage increases at a sublinear rate. This corresponds to the different curvature in the cumulative infected curves for the two scenarios up to around 70% infected. Infections seeded on the branched edge are also subject to greater variation in the final damage to the tissue between iterations of the model, especially when $F^{\text{\tiny thresh}}$ takes a value around 0.2. Finally, we also note that while the mean $F^{\infty}$ and the $F^{\infty}$ of the mean trajectory agree well for the open edge source case, where the overall variation in final cumulative damage is minimal, there is a notable difference between the two for the branched edge source case. Here, especially for $F^{\text{\tiny thresh}}\in[0.2,0.4]$, the mean trajectory sustains a substantially lower final proportion of damage compared to the average damage in the individual runs. This difference in $F^{\infty}$ suggests that stochastic effects in individual runs of the model have a substantial effect on the outcome of the immune response, according to our model, and that averaged, deterministic models (as others have used in spatial viral dynamics [7, 6]) may not fully capture the observed dynamics.

### v4

@@ 3.3 Tissue geometry and aspect ratio influence the fate of competing viral lineages
As a starting point, we sought to compare the lineage dynamics on tubes of varying circumferences. We did so by constructing tubes of length 500 cells and different choices of circumference $C$. As in an earlier result, we ran simulations of our model by seeding infection with four initially infected cells randomly placed on the left edge of the tube, with the exception that these cells are now assumed to be infected by different viral lineages. As a control, we also compared each case with a “naive” approach, where we constructed an approximately square tissue with the same number of cells, and applied toroidal periodic boundary conditions. Note that we define a tissue as “square” by having equally many cells in width as in length, however, due to the hexagonal packing of the cells, the physical dimensions of the sheet are rectangular. In the absence of tissue edges in this case, the four seed cells were placed randomly in the toroid. This control case is representative of the standard modelling approach in the literature where the geometry of the respiratory tract is not considered. Schematics and simulations on different geometries are presented in Figure 5. We show schematics of the tube and toroid geometries in Figure 5(a)–(b). We compare the results obtained for a narrow tube (with a circumference of 8 cells) to those obtained for a wide tube (circumference of 64 cells) — along with their corresponding toroids — in Figure 5(c)–(j).
@@
In order to quantify these observations, we ran 100 simulations on both the narrow and wide tube and their corresponding toroids and reported the distribution of proportions of the tissue infected by each lineage. We show these results as histograms in Figure 5(i)–(j). Figure 5(i)–(j) shows two key results. Firstly, the distributions for the toroids and tubes are qualitatively completely different. For the tubes, lineages are by far most likely to infect <10% of the tissue, with higher proportions markedly less likely, demonstrating dominance by a single lineage (or a small number of lineages). By contrast, for the toroids, density is centred at 20–30%, corresponding to a relatively even split of the tissue between the four lineages. This distribution is more tightly peaked for the larger toroid due to the larger cell population. This finding suggests that taking the typical approach to multicellular modelling, that is, in taking the cell sheet to be approximately square and with toroidal boundaries, is insufficient for capturing the lineage dynamics seen on tissue geometry more realistic to the respiratory tract. Figure 5(i)–(j) also show a small but not insignificant probability of a single lineage dominating the cell population (>80%) in the narrower tube. This is not observed for the wide tube.
3.4 Branching of the respiratory tract promotes loss of viral diversity within airway branches and spatial isolation of viral lineages
Having established the results in the previous section for a single tube, we next sought to explore how viral lineage dynamics were impacted by the inclusion of branching structure in the tissue geometry. To do so, we first explored how the lineage dynamics on the branching tree were impacted based on whether the infection started in the upper or lower branches. We conducted a set of 100 simulations on a branching tree where, as with the tubes, infection was initiated with four infected cells of different lineages, randomly placed on the left, open edge of the tree (left edge of Generation 1). Then, for comparison, we also carried out 100 simulations where the seed cells were randomly placed within a single branch on the right, branching edge of the tree (right edge of Generation 5). These two scenarios are chosen to represent, in an abstract way, infections of the upper and lower respiratory tract, respectively. As in Section 3.3, we again use a branching tree with maximum circumference of 64 cells and five generations of length 100 cells, such that the overall dimensions of the tissue are the same as that of the wide tube, as previously defined. Simulations on tube and branching geometries and analyses are shown in Figure 6. In Supplementary Videos S1 and S2 we show example simulations of infections on the branching geometry from the open edge or the branching edge, respectively.
@@
In Figure 6(e)–(f) we show representative snapshots of the unrolled branching tissue for both cases of the source cell positions, along with the tube snapshots from the previous section for comparison (Figure 6(c)–(d)). Figure 6(b) gives a schematic of these simulations. Figure 6(e)–(f) provide clear visual evidence for the sharp contrast between the lineage dynamics of infection on the branching tree starting in the upper or lower branches. These snapshots also suggest that the tissue is more likely to be dominated by a single lineage in the branching tree case (seeded on the branching edge) compared to the narrow tube case due to the larger overall cell population. As can be seen in Figure 6(f), by the time the infection spreads the length of the tree and reaches Generation 1, there is essentially only one lineage left but the entire upper half of the tree remaining to be infected, resulting in that lineage infecting a far greater overall proportion of the total cell population. The lineage dynamics, including crowding–driven extinction events, can be seen emerging over time in Supplementary Videos S1 and S2.
@@
As another descriptive tool for analysing the lineage dynamics of infections on varying tissue geometries, we next developed a notion of lineage “extinction” if a lineage is not found beyond some specified cut-off depth into the tissue (we use a cut-off of 300 cells). We denote by $P_{\text{\tiny ext}}$ the probability that, in a given geometry, a given lineage will go extinct. Figure 6(b) shows a schematic of this construction. Clearly, extinction as we have described it relies on a single direction of viral invasion, and is therefore not well defined for infection of the branching tube from the branched edge. In the narrow tube, we computed $P_{\text{\tiny ext}}=0.633$, meaning that a given lineage was more likely than not to go extinct, whereas for the wide tube, we found $P_{\text{\tiny ext}}=0.245$, meaning extinction only occurred around 25% of the time. Interestingly, on the branching tree (with source on the open edge), the probability of extinction was slightly reduced at $P_{\text{\tiny ext}}=0.200$, despite containing the same number of cells as the wide tube but with the inclusion of narrow branches within its geometry. This finding results from the fact that the branching structure acts to keep lineages separate, which slightly reduces the chance of their being crowded out. This can be seen in the lineage structure at the right end of both the wide tube and the branching tube (with source on the open edge). While in the tube case the remaining lineages are well mixed across the tube’s circumference, in the branching case, each of the narrow terminal branches is almost entirely uniform in lineage. When infection is seeded in these narrow branches, however, the branching structure does not act to separate and protect lineages since, as was observed in the narrow tube (Figure 6(c)), lineages are rapidly crowded out before they have a chance to be separated and lineage diversity is quickly lost. In brief, within branching tissues, there is a tradeoff between spatial restriction imposed by the narrowing of the branches, which promotes lineage extinction, and the physical separation of different branches, which prevents the populations of the branches of a given generation from competing with each other, thereby reducing the chance of lineage extinction. We found here that the overall effect of the branching structure is to slightly decrease lineage extinction overall compared to a tissue of the same size without branching (the wide tube geometry).
@@
To further quantify the difference in lineage structure on the branching tree (seeded on the open edge) and the wide tube, we analysed properties of the lineage dynamics along the depth of the tissue. To do so, for a given depth $d$ cells along the tissue, we examined the band of cells with depth $d-b/2$ to $d+b/2$ (where $b$ is the bandwidth, taken to be 10 cells). We sketch this computation process in Figure 7(a). For each band of cells analysed, we computed both the number of lineages found in the band, and a clustering metric, $\kappa_{\text{\tiny lin}}$, as a measure of the extent to which lineages are clustered together. This metric, which is similar to one which we developed elsewhere [32], is computed on a cell population after an infection is complete and simply reports the mean proportion of a cell’s neighbours which are of the same lineage as the cell. We show a schematic of this calculation in Figure 7(b). We plot both the number of lineages and the clustering metric $\kappa_{\text{\tiny lin}}$ as a function of band depth $d$ in Figure 7(c)–(d). In both cases we indicate the mean along with the 10th and 90th percentiles. Figure 7(c)–(d) both show that for the first approximately 200 cells of depth, the two geometries are relatively indistinguishable, and the difference in the final number of lineages remaining is not substantial compared to the noise of the system. However, for $\kappa_{\text{\tiny lin}}$, we observe that, as well as there being a notable divergence in the means for the branching tree and tube case, there is also a substantial difference in the amount of variation in each case. While $\kappa_{\text{\tiny lin}}$ is an extremely noisy metric towards the right end of the tissue, it is very consistently close to 1 for the branching tube. This confirms our observation from the snapshots that while lineages are somewhat mixed in the tube case, the branches in Generation 5 of the branching geometry generally contain only a single lineage. Note that the reason for the small range of $\kappa_{\text{\tiny lin}}$ values (around [0.7,1]) is a result of the high proportion of infections arising from cell–to–cell infection, leading to a high probability of same-lineage neighbour cells. For a different ratio of cell–to–cell to cell-free infections, it is possible that the difference in $\kappa_{\text{\tiny lin}}$ trajectories for the branching tree and tube geometries would be more pronounced.
@@ 3.5 Immune response depends on the location of infection within a branching structure
Our simulations so far have not explicitly considered the role of the immune response. Resolution of infection in our model is controlled only by the availability of the susceptible cell population, a defining feature of target cell–limited models. However, actual respiratory infections in vivo do not result in the destruction of the entire respiratory epithelium. Instead, only reasonably localised regions are affected, such as the upper airways or the alveolar passages, before the host immune response or drug intervention results in the clearing of infection [15]. We sought to investigate the behaviour of a basic model of immune activity, and how it might interact with infections seeded in different regions of the branching tree geometry.
@@
Figure 8: (a) Schematic of a simple model for the immune response in the respiratory tract. We first specify a threshold, $F^{\text{\tiny thresh}}$ for cumulative infected cell proportion. Once the infection meets the threshold, we assume the immune response is triggered, and then takes $t_{\text{act}}$ hours to activate. Once activated, we assume that the immune response immediately clears the infection. We report the cumulative infected proportion of the tissue at this point, $F^{\infty}$, as a measure of damage to the tissue. (b) Time series for the cumulative infected proportion for infections on the branching tree, where infection is initiated with a randomly placed single cell on either the left edge of Gen. 1 (the open edge of the geometry), or on the right edge of Gen. 5 (the branched edge). In each case, we plot ten trajectories and the mean, along with an indicative value of $F^{\text{\tiny thresh}}$. (c)–(d) Cumulative tissue damage on the branching tree after immune action for varying $F^{\text{\tiny thresh}}$ and fixed $t_{\text{\tiny act}}=30$h, where the infection is seeded on the open edge or the branched edge, respectively. We show violin plots for $F^{\text{\tiny thresh}}$ across 100 simulations (mean in black). We also plot in red the predicted final cumulative damage where the immune response is applied to the mean of the trajectories.
@@
As in the previous result, we sought to apply our simple model of the immune response to infections of the branching tree beginning from either the open left edge of Generation 1, or from the highly branched right edge of Generation 5. We recall that these two scenarios are chosen to represent upper– and lower–respiratory infections in vivo, respectively. In each scenario, as in Figure 3, we initiate infection with a single infected cell randomly placed on the specified edge of the tissue. We plot representative behaviour of both infection scenarios in Figure 8(b). Figure 8(b) shows the cumulative infected proportion over time for infections seeded on either the open or the branched edge. In each case we plot ten representative trajectories and the mean trajectory. The dynamics of infections progressing from either end of the tissue geometry are substantially different from each other. For infections initiated on the open edge, the cumulative infected proportion grows roughly linearly throughout, whereas for infections spreading from the branched edge, initial growth is slow and exponential, followed by a change in convexity at a cumulative infected proportion of around 0.7. Figure 8(b) shows that the trajectories of individual simulations are qualitatively similar whether the source is on the open or branched edge of the tissue.
@@
We began by testing the final cumulative damage to the tissue, $F^{\infty}$, for each of these infection seed positions for varying values of the detection threshold, $F^{\text{\tiny thresh}}$. We ran 100 simulations of the model for infections of the branching tree seeded either on the open or the branched edge, and applied our immune model for a range of $F^{\text{\tiny thresh}}$ values between 0.1 and 0.5. Here, we have kept $t_{\text{\tiny act}}$ fixed at 30h. We constructed violin plots for the distribution of $F^{\infty}$ values at each $F^{\text{\tiny thresh}}$ value, and annotated these with the mean $F^{\infty}$ value. We compared this to the $F^{\infty}$ computed on the mean trajectory for both source positions, which we also show on the same axes. That is, we compare the mean of the $F^{\infty}$ on the individual trajectories and compare this to the $F^{\infty}$ on the mean trajectory. We show these results for both the open and branched edge source in Figure 8(c) and (d), respectively. Figure 8(c)–(d) show that cumulative damage increases roughly linearly on this range of detection thresholds for the open edge case, while for the branched edge case, damage increases at a sublinear rate. This corresponds to the different curvature in the cumulative infected curves for the two scenarios up to around 70% infected. Infections seeded on the branched edge are also subject to greater variation in the final damage to the tissue between iterations of the model, especially when $F^{\text{\tiny thresh}}$ takes a value around 0.2. Finally, we also note that while the mean $F^{\infty}$ and the $F^{\infty}$ of the mean trajectory agree well for the open edge source case, where the overall variation in final cumulative damage is minimal, there is a notable difference between the two for the branched edge source case. Here, especially for $F^{\text{\tiny thresh}}\in[0.2,0.4]$, the mean trajectory sustains a substantially lower final proportion of damage compared to the average damage in the individual runs (the $F^{\infty}$ computed on the mean trajectory has a $z$-score of -1.86 compared to the $F^{\infty}$ on individual trajectories at $F^{\text{\tiny thresh}}=0.3$ and -1.56 at $F^{\text{\tiny thresh}}=0.35$; for infections seeded on the open edge of the tissue, $z$-scores do not exceed 0.14 in magnitude across the range of $F^{\infty}$ values tested). This difference in $F^{\infty}$ suggests that stochastic effects in individual runs of the model, at least for some initial conditions or perhaps in some tissue geometries, may have a substantial effect on the outcome of the immune response and that averaged, deterministic models (as others have used in spatial viral dynamics [7, 6]) may not fully capture the observed dynamics.